Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Wavefront sensing)
lumaca, nardecchia, taranto - 17:40 Friday 02 August 2024 (64844) Print this report
HWS-INJ measurement

A new measurement has been performed with HWS-INJ to verify the position of the WI DAS OUT after the movement done yesterday (see entry 64836).

All the actions are listed below:

- 10.29 UTC: manual unlock.

- 12.04 UTC: ITF in WI Single Bounce. In this configuration we checked the HWS-INJ illumination (see Fig. 1).

- 12.06 UTC: HWS-INJ acquisition started (Folder 20240802T1406).

- 12.13 UTC: the waveplate has been rotated to set the power of the WI DAS OUT up to about 4 W impinging on the CP (after WF 7).

- 13.55 UTC: HWS-INJ measurement interrupted (WF 105).

- 14.01 UTC: we restored the WI CO2 actuators powers.

Three different images of the WI DAS OUT OPL in three different moment are shown in Fig. 2.

There was still a misalignment in the horizontal direction. Thus, in agreement with Diego, in order to compensate for this misalignment, we moved BS_REC by 1cm (5556 steps) + 0.5 cm (2778 steps) in TY forward. We followed the ring movement looking at the thermal camera and the black paper of the diagnostic system (see Fig.3).
The operation has been completed at 14.12 UTC.
We left all the CO2 actuators in the 'run configuration'.

Taking advantage of the unlock at 15.34 UTC, Andrea switched the WI picomotors OFF.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×